A boat takes 3 hours to go 6 miles upstream. It can go 18 miles downstream in the same time. Find the rate of the current and the rate of the boat in still water. (Hint: Because the current pushes the boat when it is going downstream, the rate of the boat downstream is the sum of the rate of the boat and the rate of the current. The current slows down the boat when it is going upstream, so the rate of the boat upstream is th difference of the rate of the boat and the rate of the current.
Answer by stanbon(75887)

A boat takes 3 hours to go 6 miles upstream. It can go 18 miles downstream in the same time. Find the rate of the current and the rate of the boat in still water. (Hint: Because the current pushes the boat when it is going downstream, the rate of the boat downstream is the sum of the rate of the boat and the rate of the current. The current slows down the boat when it is going upstream, so the rate of the boat upstream is th difference of the rate of the boat and the rate of the current.
-----
Upstream DATA:
distance = 6 miles ; time = 3 hrs ; rate = d/t = 6/3 = 2 mph
-------
Downstream DATA:
distance = 18 miles ; time = 3 hrs ; rate = d/t = 18/3 = 6 mph
--------
Upstream rate:: b - c = 2 mph
Downstream rate:b + c = 6 mph
-------------------------------
Add to get:: 2b = 8 mph
b = 4 mph (boat speed in still water)
b + c = 6 mph
4 + c = 6 mph
c = 2 mph (speed of the current)
------------
Cheers,
Stan H.
------------